OpenStudy (anonymous):

Find the dimensions of a closed right circular cylindrical can of smallest surface area whose volume is 128pi cm^3

OpenStudy (anonymous):

I can solve it with calculus.

OpenStudy (anonymous):

so the equation you want to minimize is: \[2\pi r^2 + 2\pi rh \] you are given the condition that: \[\pi hr^2 = 128 \pi\]

OpenStudy (anonymous):

you can solve the second equation for h by dividing both sides by r^2 and dividing by pi

OpenStudy (anonymous):

the equation you want to minimize now has one variable: \[2\pi r^2 + 256\pi /r\]

OpenStudy (anonymous):

the definition of a minimum is where the slope is 0, or where the derivative is 0

OpenStudy (anonymous):

you can now take the derivative of this equation. The result is: \[4\pi r - \frac{256\pi}{r^2} \]

OpenStudy (anonymous):

set this equal to 0 and you finally have an equation that you can solve!

OpenStudy (anonymous):

factor out 4pi to get: \[4\pi (r - 64/r^2)\]

OpenStudy (anonymous):

Where did you get the 256π/r

OpenStudy (anonymous):

find a common denominator: \[4\pi (\frac{r^3 - 64}{r}) = 0\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

backtrack a little bit

OpenStudy (anonymous):

remember the first equation that we solved for h with?

OpenStudy (anonymous):

Ok I see. But, how did you get the condition as pi hr^2?

OpenStudy (anonymous):

that is the formula for volume, and the volume has to equal 128pi

OpenStudy (anonymous):

h is the height and is the radius if it wasn't obvious

OpenStudy (anonymous):

wow, yeah i'm a bit tired. So, from the simplified equation I would find the min by setting the derivative equal to zero, and checking that by taking the second derivative, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you are welcome